OpenStudy (anonymous):

Your holding a glass of lemonade out in the sun and its not so cold and refreshing anymore so you decide to put two ice cubes in the glass. Your glass of lemonande contains 0.24kg lemonade that is 33 degrees celcius, one ice cube weighs 0.025kg and are about 0 degrees celsius. If u assume that the heat loss in your enviroment is small, what is the final degree? What would the degree be with six ice cubes instead? Need help! How to approach this!

OpenStudy (vincent-lyon.fr):

Zero change is enthalpy between the initial and final states.

OpenStudy (shamim):

Heat loss=heat absorb

OpenStudy (shamim):

Use formula Q=mL Q=ms*(theta2-theta1)

OpenStudy (shamim):

Response plz!!!

OpenStudy (anonymous):

@shamim Show me more?

OpenStudy (ipwnbunnies):

I have a good idea of what to do, but I don't wanna tell you something wrong. Shamim and Vincent have the right idea, but I'm not too sure on how to set up the equation lol. Heat = mass*specific heat capacity*change in temperature So, you have to do a conservation of energy thing with this system.

OpenStudy (anonymous):

cant get it right... need to kow how to setup the equations..:(

OpenStudy (anonymous):

@shamim or @vincent-lyon.fr tell me more?

OpenStudy (shamim):

Let the final temperature is =theta

OpenStudy (shamim):

So heat losses by hot water is=m*s*(theta2-theta1)=0.24*4200*(33-theta) Joule

OpenStudy (shamim):

Now i wanna calculate heat gains by ice

OpenStudy (shamim):

Heat gains by ice to become water with 0 degree celcius=mLf=0.025*2*Lf=?

OpenStudy (shamim):

Now heat gains by ice to become theta degree celcius=ms(theta-0)=2*0.025*4200*theta

OpenStudy (shamim):

Total heat gain=?

OpenStudy (shamim):

Use Heat gain=heat loss Theta=?

OpenStudy (shamim):

Feel free to ask for more clarification!!

OpenStudy (anonymous):

@shamim what is Lf`?

OpenStudy (shamim):

Lf=latent heat of fusion

OpenStudy (anonymous):

so 333.5?

OpenStudy (shamim):

Do u know the value of Lf

OpenStudy (anonymous):

No?

OpenStudy (shamim):

Lf=334*10^3 joule/kg

OpenStudy (shamim):

Its a constant

OpenStudy (anonymous):

yea, its just not called Lf in my country :) But what do I do with: "Now heat gains by ice to become theta degree celcius=ms(theta-0)=2*0.025*4200*theta"?

OpenStudy (anonymous):

why 4200 there?

OpenStudy (shamim):

We hv to calculate total heat gain by ice

OpenStudy (anonymous):

how does the final equation look like?

OpenStudy (shamim):

U hv to add 2 heat gain. U will get total heat gain by ice

OpenStudy (anonymous):

so I put the two heat gain equations equal first? and then heat gain = heat loss?

OpenStudy (shamim):

Heat gain=2*0.025*334*10^3+2*0.025*4200*theta

OpenStudy (anonymous):

aa I see!

OpenStudy (shamim):

Ya heat gain=heat loss

OpenStudy (shamim):

Theta=?

OpenStudy (shamim):

Medal?!!!

OpenStudy (anonymous):

13.53?

OpenStudy (anonymous):

degrees celsius

OpenStudy (anonymous):

@shamim?

OpenStudy (anonymous):

but if I put 6 ice cubes? with this formula it says -10.43 degrees @shamim

OpenStudy (anonymous):

but it cant because water and ice can only reach zero?

OpenStudy (shamim):

Now i wanna talk abt 6 cubes

OpenStudy (shamim):

Can u tell me how much heat loss is happening frm water of 33 degree celcius to 0 degree celcius?

OpenStudy (shamim):

Actually this amount of heat is not enough to melt all the ice of 6 cubes

OpenStudy (shamim):

A part of ice will b melt with 0 degree celcius water and a part of ice will not melt. And this will b result

OpenStudy (shamim):

Response plz!!!